Assume the following: You are a contractor and have just been awarded a Greenfield Project. A budget has been established to install 20,000ft of pipe at an estimated 15,000 man hours and an estimated billable cost of $700,000. The original schedule planned for 4,000ft of pipe to be installed per day by a 300-man crew working a 10-hour shift (anticipated). The job was scheduled to start in Monday and finish on Friday. On Monday 2,000 ft of pipe was installed. On Tuesday 3,500 ft of pipe was installed and on Wednesday 1,000ft of pipe was installed. Actual cumulative man hours incurred to-date is 9,500 hrs. In addition, your crew has installed 1,200ft of pipe and has incurred 2,040 of man hours to-date working on Change Orders. Note: Assume that $200,000 has been billed to the client to-date.

  1. What is the current schedule variance?
  2. What is the current cost variance?
  3. How many man hours are left to be earned?
  4. How many actual man hours will be incurred in order to earn the remaining man hours?
  5. What is the current projected EAC (Estimate at Completion) in terms of man hours?
  6. What is the impact of the Change Orders on your project schedule?
  7. How much construction cost is being floated by the contractor?
  8. What are your suggestions based on the calculated results?
Answer #1
1.Current schedule variance
Upto Wednesday
Budgeted man-hrs. 300*10*3= 9000
Actual man-hrs.(9500-2040) 7460
Schedule variance= 1540
2.Current cost variance
Budgete costs upto Wednesday 700000/5= 140000
Actual billed upto Wednesday 200000
Cost Variance -60000
3.Man hours left to be earned
Budgeted man-hrs 15000
Actual man-hrs.so far(9500-2040) 7460
Man-hrs. left to be billed 7540
4.Current projected EAC (Estimate at Completion) in terms of man hours
700000/15000*7540=
351867
5.Impact of the Change Orders on your project schedule
in $ terms
700000/15000*2040=
95200
6.Construction cost being floated by the contractor
200000/(9500-2040)*7540=
202145
7.Suggestions based on the calculated results
1. Change orders have consumed almost 14% of the billable amt.estimated
2. Even with that, amt. as per 6+the impact of change orders works out to a total os $ 297345 ,ie.< the estimate as per 5.
3. The cost variance is due to the change order costs , net of which it looks favourable variance , from the estimate.
4. It is suggetsed to avoid repeat/change work.